Find the Range, Standard Deviation, and Variance, for the given data samples. 1. 93, 67, 49, 55, 92, 87, 77, 66, 73, 96, 54 2. 48, 91, 87, 93, 59, 68, 92, 100, 81 3. 8, 6, 8, 6, 8, 6, 8, 6, 8, 6, 8, 6, 8 4. -8, -5, -12, -1, 4, 7, 11 5. -23, -17, -19, -5, -4, -11, -31


1
Expert's answer
2021-12-10T04:30:53-0500

1.

"49, 54, 55, 66, 67, 73, 77, 87, 92, 93, 96"

"n=11"


"Range=96-49=47"


"mean=\dfrac{1}{n}\sum _i x_i=\dfrac{1}{11}(49+54+55+66+67""+73+77+87+92+93+96)=\dfrac{809}{11}"

"Var(X)=s^2=\dfrac{1}{n-1}\sum (x_i-\bar{x})^2"

"=\dfrac{1}{11-1}((49-\dfrac{809}{11})^2+(54-\dfrac{809}{11})^2"

"+(55-\dfrac{809}{11})^2+(66-\dfrac{809}{11})^2+(67-\dfrac{809}{11})^2"

"+(73-\dfrac{809}{11})^2+(77-\dfrac{809}{11})^2+(87-\dfrac{809}{11})^2"

"+(92-\dfrac{809}{11})^2+(93-\dfrac{809}{11})^2+(96-\dfrac{809}{11})^2)"

"\approx284.472727"

"s=\sqrt{s^2}\approx\sqrt{284.472727}\approx16.866319"

2.

"48, 59, 68, 81, 87, 91, 92, 93, 100"

"n=9"


"Range=100-48=52"


"mean=\dfrac{1}{n}\sum _i x_i=\dfrac{1}{9}(48+59+68+81+87""+91+92+93+100)=\dfrac{719}{9}"

"Var(X)=s^2=\dfrac{1}{n-1}\sum (x_i-\bar{x})^2"

"=\dfrac{1}{9-1}((48-\dfrac{719}{9})^2+(59-\dfrac{719}{9})^2"

"+(68-\dfrac{719}{9})^2+(81-\dfrac{719}{9})^2+(87-\dfrac{719}{9})^2"

"+(91-\dfrac{719}{9})^2+(92-\dfrac{719}{9})^2+(93-\dfrac{719}{9})^2)"

"+(100-\\dfrac{719}{9})^2)\\approx \t311.611111"

"s=\\sqrt{s^2}\\approx\\sqrt{ \t311.611111}\\approx17.652510"

3.

"6, 6, 6, 6, 6, 6, 8, 8, 8, 8, 8, 8, 8"

"n=13"


"Range=8-6=2"


"mean=\dfrac{1}{n}\sum _i x_i=\dfrac{1}{13}(6+6+6+6+6""+6+8+8+8+8+8+8+8)=\dfrac{92}{13}"

"Var(X)=s^2=\dfrac{1}{n-1}\sum (x_i-\bar{x})^2"

"=\dfrac{1}{13-1}((6-\dfrac{92}{13})^2+(6-\dfrac{92}{13})^2"

"+(6-\dfrac{92}{13})^2+(6-\dfrac{92}{13})^2+(6-\dfrac{92}{13})^2"

"+(6-\dfrac{92}{13})^2+(8-\dfrac{92}{13})^2+(8-\dfrac{92}{13})^2"

"+(8-\dfrac{92}{13})^2+(8-\dfrac{92}{13})^2+(8-\dfrac{92}{13})^2"

"+(8-\dfrac{92}{13})^2+(8-\dfrac{92}{13})^2\approx 1.076923"

"s=\sqrt{s^2}\approx\sqrt{1.076923}\approx1.037749"

4.

"-12, -8, -5, -1, 4, 7, 11"

"n=7"


"Range=11-(-12)=23"


"mean=\dfrac{1}{n}\sum _i x_i=\dfrac{1}{7}(-12-8-5-1+4""Var(X)=s^2=\dfrac{1}{n-1}\sum (x_i-\bar{x})^2"

"=\dfrac{1}{7-1}((-12+\dfrac{4}{7})^2+(-8+\dfrac{4}{7})^2"

"+(-5+\dfrac{4}{7})^2+(-1+\dfrac{4}{7})^2+(4+\dfrac{4}{7})^2"

"+(7+\dfrac{4}{7})^2+(11+\dfrac{4}{7})^2)\approx 69.619048"

"s=\sqrt{s^2}\approx\sqrt{69.619048}\approx8.343803"

5.

"-31, -23, -19, -17, -11, -5, -4"

"n=7"


"Range=-4-(-31)=27"


"mean=\dfrac{1}{n}\sum _i x_i=\dfrac{1}{7}(-31-23-19-17-11""-5-4)=-\dfrac{110}{7}"

"Var(X)=s^2=\dfrac{1}{n-1}\sum (x_i-\bar{x})^2"

"=\dfrac{1}{7-1}((-31+\dfrac{110}{7})^2+(-23+\dfrac{110}{7})^2"

"+(-19+\dfrac{110}{7})^2+(-17+\dfrac{110}{7})^2+(-11+\dfrac{110}{7})^2"

"+(-5+\\dfrac{110}{7})^2+(-4+\\dfrac{110}{7})^2)\\approx \t95.571429"

"s=\\sqrt{s^2}\\approx\\sqrt{\t95.571429}\\approx\t9.776064"